WHAT VOLUME OF A 60% HYDROCHLORIC ACID SOLUTION MUST BE ADDED TO A 100 ML OF A 30% HYDROCLORIC SOLUTION TO MAKE A 36% HYDROCLORIC SOLUTION?
Answer by oberobic(2304)

With mixture problems you have to keep track of how much 'pure' stuff you have or need.
.
60%*x + 30%*100 = 36% (x+100)
.60x + .30(100) = .36(x+100)
.
multiply both sides of 100
.
60x + 30(100) = 36(x+100)
60x + 3000 = 36x + 3600
24x = 600
.
x = 25
.
Answer: Add 25 ml of 60% acid to 100 ml of 30% to produce 125 ml of 36% acid.
.
.36*125 = 45 ml 'pure' acid
.6 * 25 = 15 ml 'pure' acid
.3 * 100 = 30 ml 'pure' acid
.
Correct.
.
Done.